For a party, adult tickets were sold at $15 and child tickets were sold at $10. there were 140 tickets sold for a total value of $1600. How many of each type of ticket were sold?

Answer by shweta(56)

Cost of adult ticket= $15
Cost of child ticket= $10
Number of adults= x
Number of children= y
Total number of tickets= 140
x+y = 140 .....(1)
Cost of x adult tickets= $15x
Cost of y child ticket= $10y
15x + 10y= 1600 ...(2)
x + y = 140
multiply this equation with 10 to make the coefficient of one variable same
10x+ 10y = 1400 ....(3)
Subtract equation 2 and 3
15x-10x +10y-10y= 1600-1400
5x = 200
x = 200/5
x = 40
Substitute the value of x in equation 1
x+ y =140
40 + y= 140
y= 140-40
y= 100